### v1

@@ 5.3 Spectropolarimetric observation of symbiotic star Z And
The spectropolarimetric observations of the two Raman OVI lines at and are powerful tools to diagnose the geometry of the symbiotic stars (1997A&A...321..791S; 1996A&A...310..235H; 1997A&A...327..219S). These emission lines are due to Raman scattering of the OVI resonance doublet at and by neutral hydrogen (1989A&A...211L..31S). Narrow-band Raman OVI imaging is used as a photometric tool for hunting symbiotic stars in the local Universe (2019AJ....157..156A).
Based on spectropolarimetric observations of the Raman lines, 1997A&A...327..219S determined the orbit inclination and the orbit orientation of Z And. We use this symbiotic star as a sample to explore the possibilities of FoReRo2 in the spectropolarimetric mode of observations. The symbiotic star Z And was observed on August 22, 2020 and August 10, 2021. Fig. 16 shows normalised intensity (Stokes I) and observed degree of polarisation (top panel); null parameter (middle panel); and observed position angle (bottom panel). The polarisation at and is well visible.
The null parameters are described in 2009PASP..121..993B. Both parameters are used as a quality check tool for data reduction. In Fig. 16 (middle panel) both parameters: and are close to zero at . That means the observed degree of polarisation is real.
For the wavelength calibration, one may use Ne, HgCd, and He lamps. Our experience with the spectropolarimetric data reduction shows that wavelength calibration must be applied after splitting the spectra of o- and e-beam (both for object and arc lamp) into separate fits files. Otherwise, small wavelength shifts between o- and e-beam lead to a dummy degree of polarisation in the spectral lines (for more information about small wavelength shifts and their impact on line polarisation see 2013A&A...559A.103B).
@@ 5.4 Imaging polarimetry of comet C/2019 Y4 (ATLAS)
Imaging polarimetry of comet C/2019 Y4 (ATLAS) in the wide-band filter Rx (=694.0 nm, FWHM=79.0 nm), which is chosen in such a way that the molecular emission to be avoided and only the reflected sunlight from the dust to pass through, was carried out during 7 nights in April and May 2020 in a wide range of phase angles above 55∘ during its disintegration. To select the aperture for imaging polarimetry, we use the methodology described in Subsection 3.6.
The measured degree of linear polarisation for each night is presented in Table 4. These results were fitted with the function from Eq. 11 and presented together with measurements from the comets polarimetry database (2010PDSS.8124E....K) in Figure 17.
@@
The slope at the inversion angle is computed by Eq. 12 and is h= %/deg.
| Date444Date of observations, DD/MM/YYYY | S-T-O555Phase angle (Sun-Target-Observer) in degrees | LP666Linear polarisation in % | 7771- uncertainty of LP in % |
@@ 6 Summary and conclusions
In this study, we present an overview of the 2-Channel Focal Reducer Rozhen (FoReRo2), with emphasis on its polarimetric and spectropolarimetric observing modes. To assess the instrument’s performance and stability, we analysed spectropolarimetric observations of four unpolarised standard stars over 62 epochs, and ten high-polarisation standard stars over 79 epochs. The dataset spans a decade of polarimetric and spectropolarimetric observations with FoReRo2.
The typical polarimetric accuracy of FoReRo2, estimated from high-polarisation standard stars, is better than 0.1% in the synthetic V band as well as in the red domain from imaging polarimetry.
The standard deviation on long-term observations of high polarisation standards can be used as a tool to detect the variability of the degree of polarisation. We used this methodology to demonstrate the variability of two high polarisation standard stars: HD 183143 and HD 204827.
Among the ten high-polarisation standard stars analysed in this work, HD 204827 exhibited significant short-term variability in its position angle. Based on our observations, we conclude that HD 204827 is unsuitable for use as a reliable polarimetric standard.
Based on spectropolarimetric observations, we made a statistical analysis of the Serkowski law parameters ( and ). The mean of at 0.58 with a standard deviation of . The mean of the coefficient is 1.23 with a standard deviation of 0.32.
Selection of the aperture for imaging polarimetry is performed by choosing the aperture at which the reduced Stokes parameters ( and ) converge to a well-defined value.
Only slitless spectropolarimetry gives reasonable results, while normal slit spectropolarimetry seems affected by strong instrumental polarisation. This is likely due to a slight misalignment of the 2-meter mirror relative to the optical axis appears when it was mounted after its last recoating.
@@
The symbiotic star Z And demonstrates variable polarisation in the Raman OVI line at .
The disintegrated comet ATLAS, shows maximum linear polarisation LPmax=21.9 % at phase angle of 88.6∘, similar to any ordinary comet.
To our knowledge, FoReRo2 is the only operational low-resolution optical spectropolarimeter mounted on a 2-meter-class telescope in Southeastern Europe. This work presents the first comprehensive calibration of the instrument and demonstrates its long-term stability and scientific applicability. The results provide a reliable reference baseline for future polarimetric and spectropolarimetric observations with FoReRo2. The instrument is well suited for transient phenomena, as well as for regular studies of solar system bodies, the interstellar medium, and various stellar sources.
@@ Acknowledgements.
Y.N. acknowledges partial support by grants: K-06-H58/3 ”Space weather and habitability of exoplantes: the role of coronal mass ejections and (super)flares” and K-06-M88/1 ”Interstellar medium - 3D distribution and physical characteristics”, by the Bulgarian National Science Fund. G.B., Y.N. and T.B. acknowledge partial support by grants: K-06-H88/5 “Physical properties and chemical composition of asteroids and comets - a key to increasing our knowledge of the Solar System origin and evolution.” by the Bulgarian National Science Fund. Yanko Nikolov, thanks for the hospitality of Armagh Observatory and Planetarium, and for the opportunity to work there for a period of several weeks. The research that led to these results was partially carried out with the help of infrastructure purchased under the National Roadmap for Research Infrastructure, financially coordinated by the Ministry of Education and Science of Republic of Bulgaria. Astrometry.net was used to derive the scale of the system with the new CCD cameras. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.Appendix A Imaging polarimetry of standard stars
@@
This appendix presents imaging polarimetry results for both unpolarised and polarised standard stars. Table 5 summarises the measurements for unpolarised standards, while Table 6 lists the results for polarised standard stars. The data illustrate the instrument’s performance in imaging polarimetric mode across various bands. Some of the observations were conducted using narrowband filters.
@@
Appendix B Polarised spectra of high polarisation standards.
In this appendix, we present the polarised spectra of the high polarisation standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197. The orange, red, and green dots represent the catalogue values taken from 1992AJ....104.1563S; 2021AJ....161...20P; 1983A&A...121..158S; 1990AJ.....99.1243T; 1982ApJ...262..732H, and the data presented on the Nordic Optical Telescope’s website888Source: https://www.not.iac.es/instruments/turpol/std/bd25727_note.html. The red curve represents Serkowski’s law, and the coefficients of Serkowski’s law are presented in Table LABEL:tab.Serkowski.
Appendix C 3D polarisation maps of high polarisation standards.
To demonstrate how the position angle of highly polarised standard stars is oriented, we created 3D maps. In Fig. 19, we present 3D polarisation maps in the vicinity of observed high-polarisation standards. We used data from 9286 Stars: An Agglomeration of Stellar polarisation Catalogues (2000AJ....119..923H). The colour of each star corresponds to its distance, based on data from Gaia Early Data Release 3 (2021AJ....161..147B). To match the data between the two catalogues, we used TOPCAT (2020ASPC..522...67T). The distance to BD +25 727 is unclear (presented with a blue dot in Fig. 19). High-polarisation standards are marked with a star symbol and are located at the centre of the images. The blue line of the high-polarisation standards represents catalogue values as presented in Table 1. In most cases, the position angle of the observed high-polarised standard star is aligned with the position angles of neighbouring stars. An exception to this is observed in HD 183143. The background image represents 100 m dust emission maps (1998ApJ...500..525S).

### v2

@@ 5.3 Spectropolarimetric observation of symbiotic star Z And
The spectropolarimetric observations of the two Raman OVI lines at and are powerful tools to diagnose the geometry of the symbiotic stars (Schmid and Schild (1997a); Harries and Howarth (1996); Schmid and Schild (1997b)). These emission lines are due to Raman scattering of the OVI resonance doublet at and by neutral hydrogen (Schmid (1989)). Narrow-band Raman OVI imaging is typically used as a photometric tool for hunting symbiotic stars in the Local Universe (Angeloni et al. (2019)).
Based on spectropolarimetric observations of the Raman lines, Schmid and Schild (1997b) determined the orbit inclination and the orbit orientation of Z And. We used this symbiotic star as a sample to explore the possibilities of FoReRo2 in the spectropolarimetric mode of observations. The symbiotic star Z And was observed on August 22, 2020 and August 10, 2021. In Fig. 16, we show the normalised intensity (Stokes I) and observed degree of polarisation (top panel); null parameter (middle panel); and observed PA (bottom panel). The polarisation at and is amply visible. The null parameters are described in Bagnulo et al. (2009). Both parameters were used as a quality check tool for data reduction. In Fig. 16 (middle panel) both parameters: and are close to zero at . That means the observed degree of polarisation is real.
For the wavelength calibration, we can use Ne, HgCd, and He lamps. Our experience with the spectropolarimetric data reduction shows that wavelength calibration must be applied after splitting the spectra of o- and e-beam (both for object and arc lamp) into separate fit files. Otherwise, small wavelength shifts between o- and e-beam would lead to a dummy degree of polarisation in the spectral lines (for more information about small wavelength shifts and their impact on line polarisation, see Bagnulo et al. (2013)).
@@ 5.4 Imaging polarimetry of comet C/2019 Y4 (ATLAS)
Imaging polarimetry of comet C/2019 Y4 (ATLAS) in the wide-band filter Rx (=694.0 nm, FWHM=79.0 nm), which was chosen in such a way to avoid the molecular emission and only account for the reflected sunlight from the dust to pass through. It was carried out over 7 nights in April and May 2020 in a wide range of phase angles above 55∘ during its disintegration. To select the aperture for imaging polarimetry, we used the methodology described in Section 3.7.
The measured degrees of linear polarisation for each night are presented in Table 4. These results were fitted with the function from Eq. 11 and presented together with measurements from the comets polarimetry database (Kiselev et al., 2010) in Fig. 17. This function is expressed as
@@
The slope at the inversion angle is computed via Eq. 12 and gives h= %/deg, following
| Date555Date of observations, DD/MM/YYYY | S-T-O666Phase angle (Sun-Target-Observer) in degrees | LP777Linear polarisation in % | 8881- uncertainty of LP in % |
@@ 6 Summary and conclusions
In this study, we present an overview of the 2-Channel Focal Reducer Rozhen (FoReRo2), with an emphasis on its polarimetric and spectropolarimetric observing modes. To assess the instrument’s performance and stability, we analysed spectropolarimetric observations of four unpolarised standard stars over 62 epochs, along with ten high-polarisation standard stars over 79 epochs. The dataset spans a decade of polarimetric and spectropolarimetric observations with FoReRo2.
The typical polarimetric accuracy of FoReRo2, estimated from high-polarisation standard stars, is better than 0.1% in the synthetic V band and in the red domain from imaging polarimetry.
The standard deviation on long-term observations of high polarisation standards can be used as a tool to detect the variability of the degree of polarisation. We used this methodology to demonstrate the variability of two high-polarisation standard stars: HD 183143 and HD 204827.
Among the ten high-polarisation standard stars analysed in this work, HD 204827 exhibited significant short-term variability in its PA. Based on our observations, we conclude that HD 204827 is unsuitable for use as a reliable polarimetric standard.
Based on spectropolarimetric observations, we made a statistical analysis of the Serkowski law parameters ( and ). We obtained the mean of as 0.58 with a standard deviation of . The mean of the coefficient is 1.23 with a standard deviation of 0.32.
The selection of the aperture for imaging polarimetry was done by choosing the aperture where the reduced Stokes parameters ( and ) converge to a well-defined value.
Only slitless spectropolarimetry gives reasonable results, while normal slit spectropolarimetry seems to be affected by strong instrumental polarisation. This is likely due to a slight misalignment of the 2-meter mirror relative to the optical axis, which appeared when it was mounted after its last re-coating.
@@
Symbiotic star Z And demonstrates variable polarisation in the Raman OVI line at .
Disintegrated comet ATLAS shows maximum linear polarisation LPmax=21.9 % at a phase angle of 88.6∘, similarly to any ordinary comet.
To our knowledge, FoReRo2 is the only operational low-resolution optical spectropolarimeter mounted on a 2-meter-class telescope in Southeastern Europe. This work presents the first comprehensive calibration of the instrument and demonstrates its long-term stability and scientific applicability. The results provide a reliable reference baseline for future polarimetric and spectropolarimetric observations with FoReRo2. The instrument is well-suited for transient phenomena, as well as for regular studies of solar system bodies, the interstellar medium, and various stellar sources.
@@ Acknowledgements.
Y.N. acknowledges partial support by grants: K-06-H58/3 ”Space weather and habitability of exoplanets: the role of coronal mass ejections and (super)flares” and K-06-M88/1 ”Interstellar medium - 3D distribution and physical characteristics”, by the Bulgarian National Science Fund. G.B., Y.N. and T.B. acknowledge partial support by grants: K-06-H88/5 “Physical properties and chemical composition of asteroids and comets - a key to increasing our knowledge of the Solar System origin and evolution.” by the Bulgarian National Science Fund. Yanko Nikolov, thanks for the hospitality of Armagh Observatory and Planetarium, and for the opportunity to work there for a period of several weeks. The research that led to these results was partially carried out with the help of infrastructure purchased under the National Roadmap for Research Infrastructure, financially coordinated by the Ministry of Education and Science of Republic of Bulgaria. Astrometry.net was used to derive the scale of the system with the new CCD cameras. This research has made use of NASA’s Astrophysics Data System Bibliographic Services.References
- RAMSES II: RAMan Search for Extragalactic Symbiotic Stars Project Concept, Commissioning, and Early Results from the Science Verification Phase. AJ 157 (4), pp. 156. External Links: Document, 1903.03225, ADS entry Cited by: §5.3.
- Polarimetry of small bodies of the solar system with large telescopes. J. Quant. Spec. Radiat. Transf. 112, pp. 2059–2067. External Links: Document, ADS entry Cited by: §3.7.
- The importance of non-photon noise in stellar spectropolarimetry. The spurious detection of a non-existing magnetic field in the A0 supergiant HD 92207. A&A 559, pp. A103. External Links: Document, 1309.2158, ADS entry Cited by: §5.3.
- Stellar Spectropolarimetry with Retarder Waveplate and Beam Splitter Devices. PASP 121 (883), pp. 993. External Links: Document, ADS entry Cited by: §3.1, §3.2, §5.3.
- Large Interstellar Polarisation Survey (LIPS). I. FORS2 spectropolarimetry in the Southern Hemisphere. A&A 608, pp. A146. External Links: Document, 1710.02439, ADS entry Cited by: Figure 15, §5.1.2.
- Estimating Distances from Parallaxes. V. Geometric and Photogeometric Distances to 1.47 Billion Stars in Gaia Early Data Release 3. AJ 161 (3), pp. 147. External Links: Document, 2012.05220, ADS entry Cited by: Figure 19, Appendix C.
- The RoboPol sample of optical polarimetric standards. A&A 677, pp. A144. External Links: Document, 2307.06151, ADS entry Cited by: §1.
- Spectropolarimetry of Galactic stars with anomalous extinction sightlines. A&A 615, pp. A42. External Links: Document, 1803.08043, ADS entry Cited by: Figure 15, §5.1.2, §5.1.2.
- Linear spectropolarimetry of polarimetric standard stars with VLT/FORS2. MNRAS 464 (4), pp. 4146–4159. External Links: Document, 1610.00722, ADS entry Cited by: Figure 15, §5.1.2, §5.1.2, §5.1.2.
- Polarization position angle standard stars: a reassessment of and its variability for seventeen stars based on a decade of observations. MNRAS 535 (2), pp. 1586–1615. External Links: Document, 2410.23466, ADS entry Cited by: §1.
- The Calern Asteroid Polarimetric Survey using the Torino polarimeter: assessment of instrument performances and first scientific results. MNRAS 465 (4), pp. 4335–4347. External Links: Document, ADS entry Cited by: §1.
- Wavelength dependence of position angle in polarization standards.. PASP 98, pp. 792–802. External Links: Document, ADS entry Cited by: §4.
- A novel quadruple beam imaging polarimeter and its application to Comet Tanaka-Machholz 1992 X. Ap&SS 239 (2), pp. 259–274. External Links: Document, ADS entry Cited by: §2.
- Spectropolarimetric orbits of symbiotic stars: SY Muscae.. A&A 310, pp. 235–238. External Links: ADS entry Cited by: §5.3.
- 9286 Stars: An Agglomeration of Stellar Polarization Catalogs. AJ 119 (2), pp. 923–927. External Links: Document, astro-ph/9910303, ADS entry Cited by: Figure 19, Appendix C.
- On standard polarized stars.. ApJ 262, pp. 732–739. External Links: Document, ADS entry Cited by: Appendix B, Table 1.
- Development of high resolution spectropolarimeter, LIPS. In Polarimetry in Astronomy, S. Fineschi (Ed.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4843, pp. 437–447. External Links: Document, ADS entry Cited by: §3.3.
- Exploration of the solar system with the Two-Channel Focal Reducer at the 2m-RCC telescope of Pik Terskol Observatory. Kinematika i Fizika Nebesnykh Tel Supplement 3, pp. 13–18. External Links: ADS entry Cited by: §1, §2.
- Low- and Medium-Dispersion Spectropolarimetry of Nova V475 Scuti (Nova Scuti 2003): Discovery of an Asymmetric High-Velocity Wind in a Moderately Fast Nova. AJ 132 (1), pp. 433–442. External Links: Document, astro-ph/0604341, ADS entry Cited by: §3.3.
- Instrumentation for astrophysical spectropolarimetry. chapter In Astrophysical Spectropolarimetry, J. Trujillo-Bueno, F. Moreno-Insertis, and F. Sánchez (Eds.), pp. 303–354. External Links: ADS entry Cited by: §3.4.
- Database of Comet Polarimetry, V1.0. Note: NASA Planetary Data System, id. EAR-C-COMPIL-5-DB-COMET-POLARIMETRY-V1.0 External Links: Document, ADS entry Cited by: Figure 17, §5.4.
- Wide Area Linear Optical Polarimeter North instrument I: optical design, filter design, and calibration. Journal of Astronomical Telescopes, Instruments, and Systems 10, pp. 044005. External Links: Document, 2412.00964, ADS entry Cited by: §1.
- Calibration of the polarimeter CasPol at the Nasmyth focus: Application to two extragalactic sources. Boletin de la Asociacion Argentina de Astronomia La Plata Argentina 61B, pp. 183–185. External Links: ADS entry Cited by: §1.
- Zero-polarization candidate regions for the calibration of wide-field optical polarimeters. A&A 684, pp. A132. External Links: Document, 2312.06435, ADS entry Cited by: §1.
- Transient and asymmetric dust structures in the TeV-bright nova RS Oph revealed by spectropolarimetry. A&A 679, pp. A150. External Links: Document, 2309.11288, ADS entry Cited by: §5.2.
- Spectropolarimetry of stars in the open cluster NGC 2244. In EAS2023, European Astronomical Society Annual Meeting, pp. 1422. External Links: ADS entry Cited by: Figure 15, §5.1.2.
- Interstellar polarization and extinction toward the Recurrent Nova T CrB. New A 97, pp. 101859. External Links: Document, 2201.11521, ADS entry Cited by: Figure 15, §5.1.2.
- Characterization of High-polarization Stars and Blazars with DIPOL-1 at Sierra Nevada Observatory. AJ 167 (3), pp. 137. External Links: Document, 2312.05312, ADS entry Cited by: §1.
- High-precision polarimetry of nearby stars (d ¡ 50 pc). Mapping the interstellar dust and magnetic field inside the Local Bubble. A&A 635, pp. A46. External Links: Document Cited by: §1, Table 1.
- Double Image Polarimeter—Ultra Fast: Simultaneous Three-color (BV R) Polarimeter with Electron-multiplying Charge-coupled Devices. AJ 161 (1), pp. 20. External Links: Document, ADS entry Cited by: Appendix B, §1, Table 1.
- SPARC4 - Simultaneous Polarimeter And Rapid Camera in 4 bands: first light, comissioning, and preliminary results. Bulletin of the Astronomical Society of Brazil 35, pp. 44–48. External Links: ADS entry Cited by: §1.
- Achromatic and super-achromatic zero-order waveplates. J. Quant. Spec. Radiat. Transf. 88, pp. 319–325. External Links: Document, ADS entry Cited by: §2.
- Maps of Dust Infrared Emission for Use in Estimation of Reddening and Cosmic Microwave Background Radiation Foregrounds. ApJ 500 (2), pp. 525–553. External Links: Document, astro-ph/9710327, ADS entry Cited by: Figure 19, Appendix C.
- SPHERE/ZIMPOL high resolution polarimetric imager. I. System overview, PSF parameters, coronagraphy, and polarimetry. A&A 619, pp. A9. External Links: Document, 1808.05008, ADS entry Cited by: §1.
- Spectropolarimetry of symbiotic stars: AG Draconis.. A&A 321, pp. 791–802. External Links: ADS entry Cited by: §5.3.
- The polarimetric orbit of Z Andromedae.. A&A 327, pp. 219–223. External Links: ADS entry Cited by: §5.3, §5.3.
- Identification of the emission bands at lambda lambda 6830, 7088.. A&A 211, pp. L31–L34. External Links: ADS entry Cited by: §5.3.
- The Hubble Space Telescope Northern-Hemisphere Grid of Stellar Polarimetric Standards. AJ 104, pp. 1563. External Links: Document, ADS entry Cited by: Appendix B, Figure 12, Table 1.
- New polarization measurements of HD 183143, HD 204827, and CYG OB2 Sch. No 12.. A&A 121, pp. 158–161. External Links: ADS entry Cited by: Appendix B.
- Wavelength dependence of interstellar polarization and ratio of total to selective extinction.. ApJ 196, pp. 261–290. External Links: Document, ADS entry Cited by: §5.1.1.
- Interstellar Polarization (review). In Interstellar Dust and Related Topics, J. M. Greenberg and H. C. van de Hulst (Eds.), IAU Symposium, Vol. 52, pp. 145. External Links: ADS entry Cited by: §5.1.2, §5.1.2.
- Microwave transmission. McGraw-Hill book company, inc. Cited by: §3.4.
- Dual-beam optical linear polarimetry from Southern skies: characterization of CasPol for high-precision polarimetry. Journal of Astronomical Telescopes, Instruments, and Systems 5, pp. 028002. External Links: Document, 1903.01475, ADS entry Cited by: §1, §3.1.
- IRAF in the Nineties. In Astronomical Data Analysis Software and Systems II, R. J. Hanisch, R. J. V. Brissenden, and J. Barnes (Eds.), Astronomical Society of the Pacific Conference Series, Vol. 52, pp. 173. External Links: ADS entry Cited by: §3.1.
- Astrophysical Spectropolarimetry. Cambridge University Press. External Links: ADS entry Cited by: footnote 4.
- An Atlas of Hubble Space Telescope Photometric, Spectrophotometric, and Polarimetric Calibration Objects. AJ 99, pp. 1243. External Links: Document, ADS entry Cited by: Appendix B, Table 1.
- The Imaging X-ray Polarimetry Explorer (IXPE). In Space Telescopes and Instrumentation 2016: Ultraviolet to Gamma Ray, J. A. den Herder, T. Takahashi, and M. Bautz (Eds.), Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9905, pp. 990517. External Links: Document, ADS entry Cited by: §1.
- Systematic Variations in the Wavelength Dependence of Interstellar Linear Polarization. ApJ 386, pp. 562. External Links: Document, ADS entry Cited by: Figure 15, §5.1.2, §5.1.2, §5.1.2.
- The wavelength dependence of interstellar linear polarization: stars with extreme values of lambda max.. AJ 87, pp. 695. External Links: Document, ADS entry Cited by: §5.1.2.
- The hidden symbiotic star SU Lyn - detection of flickering in U band. Bulgarian Astronomical Journal 38, pp. 83. External Links: Document, 2206.10151, ADS entry Cited by: §4.
- Continuum Foreground Polarization and Na I Absorption in Type Ia SNe. ApJ 836 (1), pp. 88. External Links: Document, 1701.05196, ADS entry Cited by: Figure 15, §5.1.2, §5.2.
Appendix A Imaging polarimetry of standard stars
@@
This appendix presents imaging polarimetry results for both unpolarised and polarised standard stars. Table 5 summarises the measurements for unpolarised standard stars, while Table 6 lists the results for polarised standard stars. The data illustrate the instrument’s performance in imaging polarimetric mode across various bands. Some observations were conducted with narrowband filters.
@@
Appendix B Polarised spectra of high-polarisation standards.
In this appendix, we present the polarised spectra of the high-polarisation standard stars: HD 7927, HD 25443, HD 161056, HD 19820, HD 204827, BD +25 727, HD 183143, BD +59 389, HD 154445, and HD 155197 (see Fig. 18). The orange, red, and green dots represent the catalogue values taken from Schmidt et al. (1992); Piirola et al. (2021); Schulz and Lenzen (1983); Turnshek et al. (1990); Hsu and Breger (1982), and the data presented on the Nordic Optical Telescope’s website999Source: https://www.not.iac.es/instruments/turpol/std/bd25727_note.html. The red curve represents Serkowski’s law, and the coefficients of Serkowski’s law are presented in Table LABEL:tab.Serkowski.
Appendix C 3D polarisation maps of high-polarisation standards.
To demonstrate how the PA of highly polarised standard stars is oriented, we created 3D maps. In Fig. 19, we present 3D polarisation maps in the vicinity of the observed high-polarisation standards. We used data from the Heiles (2000) catalogues. The colour of each star corresponds to its distance, based on data from Gaia Early Data Release 3 (Bailer-Jones et al. 2021). The distance to BD +25 727 is unclear that is why it is presented with a blue dot in Fig. 19. High-polarisation standards are marked with a star symbol and are located at the centre of the images. The blue line over the high-polarisation standards symbol represents the catalogue values of the PA and the degree of polarisation itself as presented in Table 1. In most cases, the PA of the observed high-polarised standard star is aligned with the PAs of neighbouring stars. An exception to this is observed in HD 183143. The background image represents 100 m dust emission maps (Schlegel et al. 1998).